\begin{thebibliography}{10}

\bibitem{FNM03}
Carlos Fiolhais, F.~Nogueira, and M.~Marques.
\newblock {\em A Primer in Density Functional Theory}.
\newblock Springer-Verlag, New York, 2003.

\bibitem{KS65}
W.~Kohn and L.~J. Sham.
\newblock Self-consistent equations including exchange and correlation effects.
\newblock {\em Phys. Rev.}, 140(4A):A1133--A1138, Nov 1965.

\bibitem{PGB15}
Aurora Pribram-Jones, David~A. Gross, and Kieron Burke.
\newblock {D}{F}{T}: A {T}heory {F}ull of {H}oles?
\newblock {\em Annual Review of Physical Chemistry}, 2015.

\bibitem{RG84}
Erich Runge and E.~K.~U. Gross.
\newblock Density-functional theory for time-dependent systems.
\newblock {\em Phys. Rev. Lett.}, 52(12):997, Mar 1984.

\bibitem{C96}
M.~E. Casida.
\newblock Time-dependent density functional response theory of molecular
  systems: theory, computational methods, and functionals.
\newblock In J.~M. Seminario, editor, {\em Recent developments and applications
  in density functional theory}. Elsevier, Amsterdam, 1996.

\bibitem{PGG96}
U.J.~Gossmann M.~Petersilka and E.K.U. Gross.
\newblock Excitation energies from time-dependent density-functional theory.
\newblock {\em Phys. Rev. Lett.}, 76:1212, 1996.

\bibitem{NHKFR08}
Nadine Nettelmann, Bastian Holst, Andr\'{e} Kietzmann, Martin French, Ronald
  Redmer, and David Blaschke.
\newblock Ab initio equation of state data for hydrogen, helium, and water and
  the internal structure of jupiter.
\newblock {\em The Astrophysical Journal}, 683(2):1217, 2008.

\bibitem{LHR09}
Winfried Lorenzen, Bastian Holst, and Ronald Redmer.
\newblock Demixing of hydrogen and helium at megabar pressures.
\newblock {\em Phys. Rev. Lett.}, 102:115701, Mar 2009.

\bibitem{KDLMF12}
M.~D. Knudson, M.~P. Desjarlais, R.~W. Lemke, T.~R. Mattsson, M.~French,
  N.~Nettelmann, and R.~Redmer.
\newblock Probing the interiors of the ice giants: Shock compression of water
  to 700 gpa and $3.8\text{ }\text{ }\mathbf{g}/{\mathrm{cm}}^{3}$.
\newblock {\em Phys. Rev. Lett.}, 108:091102, Feb 2012.

\bibitem{MH13}
B.~Militzer and W.~B. Hubbard.
\newblock Ab initio equation of state for hydrogen-helium mixtures with
  recalibration of the giant-planet mass-radius relation.
\newblock {\em The Astrophysical Journal}, 774(2):148, 2013.

\bibitem{WM12}
Hugh~F. Wilson and Burkhard Militzer.
\newblock Rocky core solubility in jupiter and giant exoplanets.
\newblock {\em Phys. Rev. Lett.}, 108:111101, Mar 2012.

\bibitem{WM12b}
H.~F. Wilson and B.~Militzer.
\newblock Solubility of water ice in metallic hydrogen: Consequences for core
  erosion in gas giant planets.
\newblock {\em The Astrophysical Journal}, 745(1):54, 2012.

\bibitem{MW10}
Burkhard Militzer and Hugh~F. Wilson.
\newblock New phases of water ice predicted at megabar pressures.
\newblock {\em Phys. Rev. Lett.}, 105:195701, Nov 2010.

\bibitem{WM10}
Hugh~F. Wilson and Burkhard Militzer.
\newblock Sequestration of noble gases in giant planet interiors.
\newblock {\em Phys. Rev. Lett.}, 104:121101, Mar 2010.

\bibitem{T14}
T.~Tajima.
\newblock Laser acceleration in novel media.
\newblock {\em The European Physical Journal Special Topics},
  223(6):1037--1044, 2014.

\bibitem{C13}
Daniel Clery.
\newblock Europe aims for a cut-rate superlaser to power future particle
  accelerators.
\newblock {\em Science}, 341(6147):704--705, 2013.

\bibitem{BFBKF12}
Teresa Bartal, Mark~E. Foord, Claudio Bellei, Michael~H. Key, Kirk~A. Flippo,
  Sandrine~A. Gaillard, Dustin~T. Offermann, Pravesh~K. Patel, Leonard~C.
  Jarrott, Drew~P. Higginson, Markus Roth, Anke Otten, Dominik Kraus,
  Richard~B. Stephens, Harry~S. McLean, Emilio~M. Giraldez, Mingsheng~S. Wei,
  Donald~C. Gautier, and Farhat~N. Beg.
\newblock Focusing of short-pulse high-intensity laser-accelerated proton
  beams.
\newblock {\em Nat Phys}, 8:139--142, 2012.

\bibitem{NWTZ72}
J.~Nuckolls, L.~Wood, A.~Thiessen, and G.~Zimmerman.
\newblock Laser compression of matter to super-high densities: Thermonuclear
  (ctr) application.
\newblock {\em Nature}, 239:139--142, Sept. 15 1972.

\bibitem{CFM73}
J.~S. Clarke, H.~N. Fisher, and R.~J. Mason.
\newblock Laser-driven implosion of spherical dt targets to thermonuclear burn
  conditions.
\newblock {\em Phys. Rev. Lett.}, 30:89--92, Jan 1973.

\bibitem{HK64}
P.~Hohenberg and W.~Kohn.
\newblock Inhomogeneous electron gas.
\newblock {\em Phys. Rev.}, 136(3B):B864--B871, Nov 1964.

\bibitem{CP85}
R.~Car and M.~Parrinello.
\newblock Unified approach for molecular dynamics and density-functional
  theory.
\newblock {\em Phys. Rev. Lett.}, 55(22):2471--2474, Nov 1985.

\bibitem{DKC02}
M.~P. Desjarlais, J.~D. Kress, and L.~A. Collins.
\newblock Electrical conductivity for warm, dense aluminum plasmas and liquids.
\newblock {\em Phys. Rev. E}, 66:025401, Aug 2002.

\bibitem{KCHLC10}
J.~D. Kress, James~S. Cohen, D.~A. Horner, F.~Lambert, and L.~A. Collins.
\newblock Viscosity and mutual diffusion of deuterium-tritium mixtures in the
  warm-dense-matter regime.
\newblock {\em Phys. Rev. E}, 82:036404, Sep 2010.

\bibitem{DM12}
K.~P. Driver and B.~Militzer.
\newblock All-electron path integral monte carlo simulations of warm dense
  matter: Application to water and carbon plasmas.
\newblock {\em Phys. Rev. Lett.}, 108:115502, Mar 2012.

\bibitem{M65}
N.~D. Mermin.
\newblock Thermal properties of the inhomogenous electron gas.
\newblock {\em Phys. Rev.}, 137:A: 1441, 1965.

\bibitem{SD14}
Travis Sjostrom and J\'er\^ome Daligault.
\newblock Gradient corrections to the exchange-correlation free energy.
\newblock {\em Phys. Rev. B}, 90:155109, Oct 2014.

\bibitem{KSDT14}
Valentin~V. Karasiev, Travis Sjostrom, James Dufty, and S.~B. Trickey.
\newblock Accurate homogeneous electron gas exchange-correlation free energy
  for local spin-density calculations.
\newblock {\em Phys. Rev. Lett.}, 112:076403, Feb 2014.

\bibitem{PD84}
Francois Perrot and M.~W.~C. Dharma-wardana.
\newblock Exchange and correlation potentials for electron-ion systems at
  finite temperatures.
\newblock {\em Phys. Rev. A}, 30:2619--2626, Nov 1984.

\bibitem{TI86}
Shigenori Tanaka and Setsuo Ichimaru.
\newblock Thermodynamics and correlational properties of finite-temperature
  electron liquids in the singwi-tosi-land-sjolander approximation.
\newblock {\em Journal of the Physical Society of Japan}, 55(7):2278--2289,
  1986.

\bibitem{DAC86}
R.~G. Dandrea, N.~W. Ashcroft, and A.~E. Carlsson.
\newblock Electron liquid at any degeneracy.
\newblock {\em Phys. Rev. B}, 34(4):2097--2111, Aug 1986.

\bibitem{RA00}
J.~J. Rehr and R.~C. Albers.
\newblock Theoretical approaches to x-ray absorption fine structure.
\newblock {\em Rev. Mod. Phys.}, 72:621--654, Jul 2000.

\bibitem{FBCR10}
G\'erald Faussurier, Christophe Blancard, Philippe Coss\'e, and Patrick
  Renaudin.
\newblock Equation of state, transport coefficients, and stopping power of
  dense plasmas from the average-atom model self-consistent approach for
  astrophysical and laboratory plasmas.
\newblock {\em Physics of Plasmas}, 17(5):--, 2010.

\bibitem{FBBBB13}
A.~Frank, A.~Bla\ifmmode \check{z}\else \v{z}\fi{}evi\ifmmode~\acute{c}\else
  \'{c}\fi{}, V.~Bagnoud, M.~M. Basko, M.~B\"orner, W.~Cayzac, D.~Kraus,
  T.~He\ss{}ling, D.~H.~H. Hoffmann, A.~Ortner, A.~Otten, A.~Pelka, D.~Pepler,
  D.~Schumacher, An. Tauschwitz, and M.~Roth.
\newblock Energy loss and charge transfer of argon in a laser-generated carbon
  plasma.
\newblock {\em Phys. Rev. Lett.}, 110:115001, Mar 2013.

\bibitem{ZFGLC15}
A.~B. Zylstra, J.~A. Frenje, P.~E. Grabowski, C.~K. Li, G.~W. Collins,
  P.~Fitzsimmons, S.~Glenzer, F.~Graziani, S.~B. Hansen, S.~X. Hu, M.~Gatu
  Johnson, P.~Keiter, H.~Reynolds, J.~R. Rygg, F.~H. S\'eguin, and R.~D.
  Petrasso.
\newblock Measurement of charged-particle stopping in warm dense plasma.
\newblock {\em Phys. Rev. Lett.}, 114:215002, May 2015.

\bibitem{HBNST06}
A~P Horsfield, D~R Bowler, H~Ness, C~G S\'anchez, T~N Todorov, and A~J Fisher.
\newblock The transfer of energy between electrons and ions in solids.
\newblock {\em Reports on Progress in Physics}, 69(4):1195, 2006.

\bibitem{BSCKS12}
Lorin~X. Benedict, Michael~P. Surh, John~I. Castor, Saad~A. Khairallah,
  Heather~D. Whitley, David~F. Richards, James~N. Glosli, Michael~S. Murillo,
  Christian~R. Scullard, Paul~E. Grabowski, David Michta, and Frank~R.
  Graziani.
\newblock Molecular dynamics simulations and generalized lenard-balescu
  calculations of electron-ion temperature equilibration in plasmas.
\newblock {\em Phys. Rev. E}, 86:046406, Oct 2012.

\bibitem{CVG13}
D.~A. Chapman, J.~Vorberger, and D.~O. Gericke.
\newblock Reduced coupled-mode approach to electron-ion energy relaxation.
\newblock {\em Phys. Rev. E}, 88:013102, Jul 2013.

\bibitem{VG14}
J.~Vorberger and D.O. Gericke.
\newblock Comparison of electron-ion energy transfer in dense plasmas
  obtained from numerical simulations and quantum kinetic theory.
\newblock {\em High Energy Density Physics}, 10:1 -- 8, 2014.

\bibitem{GR09}
Siegfried~H. Glenzer and Ronald Redmer.
\newblock X-ray thomson scattering in high energy density plasmas.
\newblock {\em Rev. Mod. Phys.}, 81:1625--1663, Dec 2009.

\bibitem{LL85}
T.~Li and Y.~Li.
\newblock {\em Phys. Rev. A}, 31:3970, 1985.

\bibitem{LT85}
T.~Li and P.~Tong.
\newblock {\em Phys. Rev. A}, 31:1950, 1985.

\bibitem{YMB12}
Zeng hui Yang, Neepa~T. Maitra, and Kieron Burke.
\newblock The effect of cusps in time-dependent quantum mechanics.
\newblock {\em Phys. Rev. Lett.}, 108:063003, Feb 2012.

\bibitem{L01}
Robert van Leeuwen.
\newblock Key concepts in time-dependent density-functional theory.
\newblock {\em International Journal of Modern Physics B}, 15:1969--2023, 2001.

\bibitem{RL11}
M.~Ruggenthaler and R.~van Leeuwen.
\newblock Global fixed-point proof of time-dependent density-functional theory.
\newblock {\em EPL (Europhysics Letters)}, 95(1):13001, 2011.

\bibitem{GK90}
EKU Gross and W~Kohn.
\newblock Time-dependent density functional theory.
\newblock {\em Adv. Quant. Chem}, 21:255--291, 1990.

\bibitem{MSB03}
Neepa~T. Maitra, Ivo Souza, and Kieron Burke.
\newblock Current-density functional theory of the response of solids.
\newblock {\em Phys. Rev. B}, 68(4):045109, Jul 2003.

\bibitem{KL57}
W.~Kohn and J.M. Luttinger.
\newblock Quantum theory of electrical transport phenomena.
\newblock {\em Phys. Rev.}, 108:590, 1957.

\bibitem{PPGB14}
Aurora Pribram-Jones, Stefano Pittalis, E.K.U. Gross, and Kieron Burke.
\newblock Thermal density functional theory in context.
\newblock In Frank Graziani, Michael~P. Desjarlais, Ronald Redmer, and
  Samuel~B. Trickey, editors, {\em Frontiers and Challenges in Warm Dense
  Matter}, volume~96 of {\em Lecture Notes in Computational Science and
  Engineering}, pages 25--60. Springer International Publishing, 2014.

\bibitem{CDK15}
Debajit Chakraborty, James Dufty, and Valentin~V. Karasiev.
\newblock Chapter two - system-size dependence in grand canonical and canonical
  ensembles.
\newblock In John~R. Sabin and Remigio Cabrera-Trujillo, editors, {\em Concepts
  of Mathematical Physics in Chemistry: A Tribute to Frank E. Harris - Part A},
  volume~71 of {\em Advances in Quantum Chemistry}, pages 11 -- 27. Academic
  Press, 2015.

\bibitem{SL13}
G.~Stefanucci and R.~van {L}eeuwen.
\newblock {\em {Nonequilibrium Many-Body Theory of Quantum Systems}}.
\newblock Cambridge University Press, 2013.

\bibitem{ABBW09}
Jean-Pierre Antoine, Robert~C. Bishop, Arno Bohm, and Sujeev Wickramasekara.
\newblock Rigged hilbert spaces in quantum physics.
\newblock In Daniel Greenberger, Klaus Hentschel, and Friedel Weinert, editors,
  {\em Compendium of Quantum Physics}, pages 640--650. Springer Berlin
  Heidelberg, 2009.

\bibitem{G15}
K.~J.~H. Giesbertz.
\newblock Invertibility of retarded response functions for laplace
  transformable potentials: Application to one-body reduced density matrix
  functional theory.
\newblock {\em The Journal of Chemical Physics}, 143(5), 2015.

\bibitem{PYTB14}
A.~Pribram-Jones, Z.-H. Yang, J.~R. Trail, K.~Burke, R.~J. Needs, and C.~A.
  Ullrich.
\newblock Excitations and benchmark ensemble density functional theory for two
  electrons.
\newblock {\em J. Chem. Phys.}, 140:18A541, 2014.

\bibitem{G16}
K.J.H. Giesbertz.
\newblock Invertibility of the retarded response functions for initial mixed
  states: application to one-body reduced density matrix functional theory.
\newblock {\em Molecular Physics}.
\newblock published online.

\bibitem{YB13}
Zeng-hui Yang and Kieron Burke.
\newblock Nonexistence of a taylor expansion in time due to cusps.
\newblock {\em Phys. Rev. A}, 88:042514, Oct 2013.

\bibitem{MK87}
D.~Mearns and W.~Kohn.
\newblock {\em Phys. Rev. A}, 35:4796, 1987.

\bibitem{GK85}
E.K.U. Gross and W.~Kohn.
\newblock Local density-functional theory of frequency-dependent linear
  response.
\newblock {\em Phys. Rev. Lett.}, 55:2850, 1985.

\bibitem{Yc88}
W.~Yang.
\newblock Dynamic linear response of many-electron systems: An integral
  formulation of density-functional theory.
\newblock {\em Phys. Rev. A}, 38:5512, 1988.

\bibitem{KBP94}
L.P. Kadanoff, G.~Baym, and D.~Pines.
\newblock {\em Quantum Statistical Mechanics}.
\newblock Advanced Books Classics Series. Addison-Wesley, 1994.

\bibitem{BW12}
Kieron Burke and Lucas~O. Wagner.
\newblock Dft in a nutshell.
\newblock {\em Int. J. Quant. Chem.}, 2012.

\bibitem{Liebsch97}
Ansgar Liebsch.
\newblock {\em Electronic Excitations at Metal Surfaces}.
\newblock Physics of Solids and Liquids. Springer, 1997.

\bibitem{LP77}
D.C. Langreth and J.P. Perdew.
\newblock Exchange-correlation energy of a metallic surface: Wave-vector
  analysis.
\newblock {\em Phys. Rev. B}, 15:2884, 1977.

\bibitem{FW71}
A.~L. {Fetter} and J.~D. {Walecka}.
\newblock {\em {Quantum theory of many-particle systems}}.
\newblock McGraw-Hill, New York, NY, 1971.

\bibitem{U11}
Carsten~A Ullrich.
\newblock {\em Time-dependent Density-Functional Theory: Concepts and
  Applications}.
\newblock Oxford University Press, 2011.

\bibitem{D12}
John~F. Dobson.
\newblock {\em Dispersion (van der {W}aals) {F}orces and {T}{D}{D}{F}{T}},
  pages 417--441.
\newblock Number 837 in Lecture Notes in Physics. Springer, 2012.

\bibitem{PB15}
A~Pribram-Jones and K~Burke.
\newblock Connection formula for thermal density functional theory.
\newblock {\em Phys. Rev.}, 2015.
\newblock Submitted.

\bibitem{BWG05}
Kieron Burke, Jan Werschnik, and E.~K.~U. Gross.
\newblock Time-dependent density functional theory: Past, present, and future.
\newblock {\em The Journal of Chemical Physics}, 123(6):062206, 2005.

\bibitem{EYF10}
Henk Eshuis, Julian Yarkony, and Filipp Furche.
\newblock Fast computation of molecular random phase approximation correlation
  energies using resolution of the identity and imaginary frequency
  integration.
\newblock {\em The Journal of Chemical Physics}, 132(23):234114, 2010.

\bibitem{EF11}
Henk Eshuis and Filipp Furche.
\newblock A parameter-free density functional that works for noncovalent
  interactions.
\newblock {\em The Journal of Physical Chemistry Letters}, 2(9):983--989, 2011.

\bibitem{PRRS12}
Joachim Paier, Xinguo Ren, Patrick Rinke, Gustavo~E Scuseria, Andreas Gruneis,
  Georg Kresse, and Matthias Scheffler.
\newblock Assessment of correlation energies based on the random-phase
  approximation.
\newblock {\em New Journal of Physics}, 14(4):043002, 2012.

\bibitem{PK98b}
S.~Kurth and J.~P. Perdew.
\newblock In G.~J. Kalman, J.~M. Rommel, and K.~Blagoev, editors, {\em Strongly
  Coupled Coulomb Systems}. Plenum Press, New York, NY, 1998.

\bibitem{BPW97}
Kieron Burke, John~P. Perdew, and Y.~Wang.
\newblock {\em Derivation of a generalized gradient approximation: The PW91
  density functional}, page~81.
\newblock Plenum, NY, 1997.

\bibitem{WYB12}
Lucas Wagner, Zeng hui Yang, and Kieron Burke.
\newblock {\em Exact conditions and their relevance in TDDFT}, chapter~5, pages
  101--122.
\newblock Number 837 in Lecture Notes in Physics. Springer, 2012.

\bibitem{LP85}
M.~Levy and J.P. Perdew.
\newblock Hellmann-feynman, virial, and scaling requisites for the exact
  universal density functionals. shape of the correlation potential and
  diamagnetic susceptibility for atoms.
\newblock {\em Phys. Rev. A}, 32:2010, 1985.

\bibitem{HPB99}
Paul Hessler, Jang Park, and Kieron Burke.
\newblock Several theorems in time-dependent density functional theory.
\newblock {\em Phys. Rev. Lett.}, 82(2):378--381, Jan 1999.
\newblock {\em ibid.} {\bf 83}, 5184(E) (1999).

\bibitem{PPFS11}
S.~Pittalis, C.~R. Proetto, A.~Floris, A.~Sanna, C.~Bersier, K.~Burke, and
  E.~K.~U. Gross.
\newblock Exact conditions in finite-temperature density-functional theory.
\newblock {\em Phys. Rev. Lett.}, 107:163001, Oct 2011.

\bibitem{HB08}
Maria Hellgren and Ulf von Barth.
\newblock Linear density response function within the time-dependent
  exact-exchange approximation.
\newblock {\em Phys. Rev. B}, 78:115107, Sep 2008.

\bibitem{SPB15}
Justin~C. Smith, Aurora Pribram-Jones, and Kieron Burke.
\newblock Thermal corrections to density functional simulations of warm dense
  matter, 2015.
\newblock submitted.

\end{thebibliography}

\documentclass[twocolumn,superscriptaddress,showpacs,amsmath,amssymb,prl,longbibliography,nofootinbib]{revtex4-1}

\usepackage{listings}
\usepackage{dcolumn}  
\usepackage{bm}       
\usepackage[usenames,dvipsnames]{color}
\definecolor{URLCOL}{rgb}{0,0.52,0.83} 
\definecolor{LINKCOL}{rgb}{0.05,0.5,0} 
\definecolor{orange}{rgb}{0.6,0.3,0} 
\definecolor{CITECOL}{rgb}{0.25,0,0.48} 
\usepackage{epstopdf}

\usepackage[pdftex,bookmarks,breaklinks,bookmarksopen,bookmarksnumbered,colorlinks,linkcolor=LINKCOL,linktocpage,citecolor=CITECOL,urlcolor=URLCOL,pdfpagemode=UseOutline,pdftex]{hyperref}

\usepackage{booktabs}
\usepackage{comment}

\definecolor{TITLECOL}{rgb}{0.1,0.2,0.7} 
\definecolor{SECOL}{rgb}{0.1,0.2,0.7} 
\definecolor{CONTENTSCOL}{rgb}{0.1,0.2,0.7} 
\definecolor{SSECOL}{rgb}{0.25,0,0.48} 
\definecolor{SSSECOL}{rgb}{0.2,0.08,0.53} 
\definecolor{FINCOL}{rgb}{0.01,0.3,0.07} 

\def\check#1{{\color{orange}#1}} 

\definecolor{URLCOL}{rgb}{0,0.17,0.43} 
\definecolor{LINKCOL}{rgb}{0.05,0.4,0} 
\definecolor{CITECOL}{rgb}{0.35,0,0.48} 

\def\bea{\begin{eqnarray}}
\def\eea{\end{eqnarray}}
\def\ben{\begin{equation}}

\def\een{\end{equation}}
\def\benu{\begin{enumerate}}
\def\enu{\end{enumerate}}
\def\bei{\begin{itemize}}
\def\eei{\end{itemize}}

\def\br{{\bf r}}

\def\n{n}

\def\sss{\scriptscriptstyle\rm}

\def\g{_\gamma}

\def\x{_{\sss X}}
\def\c{_{\sss C}}
\def\s{_{\sss S}}
\def\xc{_{\sss XC}}

\def\Hxc{_{\sss HXC}}

\def\H{_{\sss H}}

\def\ee{_{\rm ee}}

\def\wt{\texttt{w}}
\def\b{^\tau}

\def\t{^{\tau}}

\def\tl{^{\tau,\lambda}}
\def\g{_\gamma}

\begin{document}

\title{
Thermal density functional theory:  Time-dependent linear response and 
approximate functionals from the fluctuation-dissipation theorem
}
\author{Aurora Pribram-Jones}
\affiliation{Lawrence Livermore National Laboratory, Livermore, CA 94550}
\affiliation{Department of Chemistry, University of California, Berkeley, CA 94720}
\author{Paul E. Grabowski}
\affiliation{Department of Physics and Astronomy, University of California, Irvine, CA 92697}
\author{Kieron Burke}
\affiliation{Department of Chemistry, University of California, Irvine, CA 92697}
\affiliation{Department of Physics and Astronomy, University of California, Irvine, CA 92697}
\date{\today}

\begin{abstract}
The van Leeuwen proof of linear-response 
time-dependent density functional theory (TDDFT) is generalized to
thermal ensembles.  This allows generalization
to finite temperatures of the Gross-Kohn relation,
the exchange-correlation kernel of TDDFT, 
and fluctuation dissipation theorem for DFT.  
This produces a natural method for generating new thermal
exchange-correlation (XC) approximations.
\end{abstract}

\pacs{%
31.15.E-, 
71.15.Mb, 
05.30.Fk 
}

\maketitle

Kohn-Sham density functional theory (KS-DFT) is a popular and well-established approach
to electronic structure problems in many areas, especially materials science and chemistry\cite{FNM03}.
The Kohn-Sham method imagines a fictitious system of non-interacting fermions with the
same density as the real system\cite{KS65} and from which the ground-state energy can be extracted.
Only a small fraction of the total energy, called the exchange-correlation (XC) energy,
need be approximated to solve any ground-state electronic problem\cite{FNM03}, and modern approximations
usually produce sufficient accuracy to be useful\cite{PGB15}.
The advent of TDDFT generalized this method to time-dependent problems\cite{RG84}.  Limiting TDDFT to linear-response
yields a method for extracting electronic excitations\cite{C96,PGG96},
once another functional, the XC
kernel, is also approximated.

But there is growing interest in systems in which the electrons are not close to zero temperature.
Warm dense matter (WDM) is partially ionized, solid-density matter having a temperature near
the Fermi energy. 
It has wide-ranging applications including
the astrophysics of giant planets and white dwarf atmospheres\cite{NHKFR08,LHR09,KDLMF12,MH13,WM12,WM12b,MW10,WM10},
cheap and ultra-compact particle accelerators and
radiation sources\cite{T14,C13,BFBKF12}, and
the eventual production of clean, abundant
energy via inertial confinement fusion\cite{NWTZ72,CFM73}.
One of the most successful methods for simulating equilibrium warm dense matter combines
DFT\cite{HK64,KS65} and molecular dynamics\cite{CP85} to capture quantum mechanical effects
of WDM electrons and the classical behavior of ions\cite{NHKFR08,LHR09,KDLMF12,MH13,WM12,WM12b,MW10,WM10,DKC02,KCHLC10,DM12}.
Such simulations use the Mermin theorem\cite{M65} to generate a KS scheme at finite
temperature, defined to generate the equilibrium density and free energy.
In practice, the XC free energy is almost always approximated with a ground-state
approximation, but formulas for thermal corrections are being developed\cite{SD14,KSDT14,PD84,TI86,DAC86}.

Many processes of interest involve perturbing an equilibrium system with some
time-dependent (TD) perturbation, such as a laser field\cite{RA00} or a rapidly moving nucleus as in 
stopping power\cite{FBCR10,FBBBB13,ZFGLC15}.
Of great interest within the WDM community are calculations of spectra, dynamic structure factors, 
and the flow of energy between electrons and ions\cite{HBNST06,BSCKS12,CVG13,VG14}.  
Spectra expose a material's response to excitation by electromagnetic radiation, which would
facilitate experimental design and analysis.  Dynamic structure factors can be
related to the x-ray scattering response, which is being developed as
a temperature and structural diagnostic tool for WDM\cite{GR09}.
Thus it would appear that a TD version of the Mermin formalism is required.
A theorem is proven in Li et al.\cite{LL85,LT85}, but the formalism assumes the
temperature is fixed throughout the process, and so cannot describe e.g., equilibration
between electrons and ions.  Moreover,
the proof requires the Taylor expansion of the perturbing potential as a function
of time, just as in the Runge-Gross (RG) theorem\cite{RG84}.  This can be problematic
for initial states with cusps\cite{YMB12}, such as at the nuclear centers. (Recent efforts\cite{L01,RL11} have focused on avoiding these complications
at zero temperature.)
Finally, the RG proof requires
invocation of a boundary condition to complete the
one-to-one correspondence between
density and potential\cite{GK90}, which create subtleties when 
applied to extended systems\cite{MSB03}.

In the present work, we prove the RG
theorem at finite temperature within linear response by generalizing the
elegant linear response proof of van Leeuwen\cite{L01} to thermal ensembles.
Our proof avoids several of the
drawbacks mentioned above, while still providing a solid grounding to 
much of WDM theoretical work.  We then define the exchange-correlation kernel at
finite temperature and generalize the Gross-Kohn equation.  Finally, we extend
the fluctuation-dissipation theorem of ground-state DFT to finite temperatures,
and show how this provides a route to {\em equilibrium} free energy XC approximations.

Consider a system of electrons in thermal and particle equilibrium with a bath
at some temperature, $\tau$, and with static equilibrium density $\n\b(\br)$.
The system extends throughout space with a finite average density, i.e., the
thermodynamic limit has been taken.  The limit of isolated atoms or
molecules is achieved by then taking the separation between certain nuclei to infinity.
In this sense, no surface boundary condition need be invoked\cite{GK90},
as the density never quite vanishes, while 
the average particle number per atom or molecule molecule is finite.
These electrons are perturbed at $t=0$
by a potential $\delta v(\br,t)$ that is Laplace-transformable.
To avoid complex questions of equilibration, we consider only the
linear response of the system, so that the perturbation does not affect
the temperature of the system as, e.g., Joule heating is a higher
order effect\cite{KL57}.
The Kubo response formula for the density change in response to $\delta v$
is
\ben
\delta\n\b(\br,s) = \int d^3r'\, \chi\b(\br,\br',s)\, \delta v(\br',s),
\label{dn}
\een
where the Laplace transform
\ben
\delta v(\br,s) = \int_{0}^\infty dt\, e^{-st}\, \delta v(\br,t)
\een
is assumed to exist for all $s>0$.
Within the grand canonical ensemble\cite{PPGB14,CDK15}, the equilibrium 
density-density response function is\cite{SL13}:
\ben
\chi\b(\br,\br',s)=
i\sum_{ij}\, \wt_i \frac{\Delta n^{\tau *}_{ij}(\br)\Delta \n\t_{ij}(\br')}
{s-i\omega_{ji}} + c.c.,
\een
where
\ben
\Delta \n\t_{ij}(\br) = \langle i | \hat \n(\br) | j \rangle - \delta_{ij}\
\n\b(\br)
\label{eqn:Dn}
\een
are matrix elements of the density fluctuation operator. The energy-ordered indices $i,j$ run over all many-body states (both bound and 
continuum\cite{ABBW09}) with all particle numbers,
but $\Delta \n\t_{ij}$ vanishes unless $N_i=N_j$. The transition frequencies 
$\omega_{ji}=E_j-E_i$, and the statistical weights $\wt_i$  are thermal occupations 
for the equilibrium statistical operator 
$\hat{\Gamma}^\tau=\sum_{i}\wt_i |\Psi_i\rangle\langle \Psi_i |$ and obey
$\wt_i<\wt_j$ if $E_i>E_j$ and $N_i=N_j$. This condition is satisfied by the grand canonical
ensemble of common interest with
$\wt_i=\exp[-(E_i-\mu N_i)/\tau]/\sum_{i}\exp[-(E_i-\mu N_i)/\tau]$.

We also need the (Laplace-transformed) one-body potential
operator:
\ben
\delta \hat V(s) = \int d^3r\, \hat \n(\br)\, \delta v(\br,s),
\een
and its matrix elements:
\ben
\delta V_{ij}(s) = \langle i | \delta\hat V(s) | j \rangle.
\label{eqn:matel}
\een
Its expectation value is
\ben
\delta V\b(s) = \sum_i \wt_i\, V_{ii}(s)
= \int d^3r\, \n\b(r)\, \delta v(\br,s),
\een
so that matrix elements of its fluctuations are
\ben
\Delta V\t_{ij}(s) = \delta V_{ij}(s)- \delta_{ij}\, \delta V\b(s).
\een
Then consider the expectation value:
\ben
m\b(s) = \int d^3r\, \delta\n\b(\br,s)\, \delta v(\br,s).
\een
Inserting Eq. (\ref{dn}) and using the definitions, we find
\ben
m\b(s) = 
-\sum_{ij}\wt_i \mid\Delta V\t_{ij}(s)\mid^2
\frac{2\omega_{ji}}{s^2+\omega_{ji}^2}.
\een
This is rearranged as
\ben
\label{eqn:pos}
m\b(s)=-2\sum_{i=0}^\infty\sum_{j=i+1}^\infty
\frac{(\wt_i-\wt_j)\omega_{ji}}{s^2+\omega_{ji}^2}\mid\Delta V\t_{ij}(s)\mid^2.
\een
We have ordered all states by energy regardless of particle number here for simplicity, though this is not strictly necessary since different particle number subsystems do not interact. 
For now, we assume no degeneracies.
Then the above expression, $m\b(s)$, vanishes
only if
every 
$\Delta V\t_{ij}(s)$ does for $i\neq j$ because of our assumption that
$(\wt_i-\wt_j)\omega_{ji}>0$ if $i\ne j$. 

The usual statement of the RG theorem is that no two potentials
that differ by more than an inconsequential function of time alone can give
rise to the same density (for fixed statistics, interparticle interaction, and 
initial state\cite{RG84}).   Imagine two such perturbations exist, yielding
the same density response.  
Since, in linear response, the density response is proportional to the perturbation, we can
subtract one from the other, and the statement to be proved is that
there is no non-trivial perturbation with zero density response.
If it did exist, then $m\b(s)$ would vanish and our algebra shows that
every $\Delta V\t_{ij}(s)$ with $i\ne j$ would also. Finally,
\ben
\sum_{k=1}^{N_i}\delta v(\br_k,s)
\Psi_j(\br_1\ldots\br_{N_i})=\sum_i \delta 
V_{ij}(s) \Psi_i(\br_1\ldots\br_{N_i}),
\label{eqn:tdep}
\een
which can be proven by integrating over all coordinates with $\Psi_k^*$.
Then, as $\Delta V\t_{ij}(s) = \delta V_{ij}(s)$ for $i\neq j$,
and must vanish if there is no density response, the sum on the right of
Eq. (\ref{eqn:tdep}) collapses to just the $j$-th term, showing that
$\delta v(\br,s)$ must be spatially independent.

We can also include a finite number ($M$) of degenerate excited eigenstates.  (For the complications involved
when the ground-state is degenerate, see Ref. \cite{G15}).
For such states, $\omega_{ij}=0$ and the argument above no longer
implies $\delta V_{ij}(s)$ vanishes, as the perturbation couples degenerate states within the same subspace. But simply choose at least $M$
points in the $3N$-dimensional coordinate space that are not on any
nodal hypersurface of the degenerate subspace.  Then the only solution to Eq. (\ref{eqn:tdep})
is again that $\delta v (\br,s)$ must be independent of $\br$.
%

Thus we have generalized the
van Leeuwen proof to thermal ensembles, even with finite degeneracies
among excited states.
Our proof applies to {\em any} ensemble with 
weights that monotonically decrease with increasing energy for
each particle number\cite{PYTB14,G16}. This avoids complications caused by cusps in initial 
wavefunctions\cite{YMB12,YB13}. 
Extension to spatially periodic potentials is straightforward, as no
boundary condition\cite{GK90} was invoked\cite{MSB03}.

In order for the above result to be of practical use, we consider the KS scheme
for finite-temperature, time-dependent systems and provide a
method for generating XC approximations.
We assume the equilbrium Mermin-Kohn-Sham (MKS)\cite{M65,KS65} potential exists.  At this point,
we switch to using the more familiar Fourier-transform notation, but in fact
all results and definitions apply only to Laplace-transformable perturbations. (In practice, this distinction rarely matters, but occasional formal difficulties
arise if this restriction is not made, see Ref. \cite{MK87} and Sec. 3.2 of Ref. \cite{L01}.)
First
we generalize the Gross-Kohn response formula\cite{GK85} to thermal ensembles. Define
\begin{align}
\chi\b(\br,\br',\omega)&=\notag\\
\sum_{jk}\, \wt_j &\left\{\frac{\langle j | \hat \n(\br) | k \rangle \langle k | \hat \n(\br') | j \rangle}
{\omega-\omega_{kj}+i\eta} - \frac{\langle j | \hat \n(\br') | k \rangle \langle k | \hat \n(\br) | j \rangle}
{\omega+\omega_{kj}+i\eta}\right\},
\end{align}
where $\eta\to 0^+$\cite{Yc88}.

Because of our proof of one-to-one correspondence,
we can invert the response function (excluding a constant), and write
\ben
\left(\chi\t\right)^{-1}(12) = \frac{\delta v(1)}{\delta \n(2)},
\label{chiinv}
\een
where $1$ denotes the coordinates $\br,t$, and 2 another pair\cite{KBP94}.
The standard definition of XC is:
\ben
v\s(1) = v(1) + v\H(1) + v\xc(1),
\een
where $v\s$ is the one-body potential of the non-interacting KS system and $v\H$ is the Hartree potential\cite{BW12}. Differentiating with respect to $\n(2)$, this yields
\ben
\left(\chi\t\s\right)^{-1}(12)=\left(\chi\t\right)^{-1}(12) + f\H(12) + f\xc\t(12),
\label{kernel}
\een
which defines the XC kernel at finite temperature, where 
$\chi\s\t$ is the KS response function\cite{Yc88} and the traditionally defined Hartree
contribution is simply
\ben
f\H(12)=\frac{\delta(t_1-t_2)}{\mid \br_1-\br_2 \mid}.
\een
This follows the definition within the Mermin formalism\cite{M65}
(but see Refs. \cite{PPGB14} and \cite{PYTB14} for alternative choices and their consequences).
Inverting yields the thermal Gross-Kohn equation\cite{GK85}:
\ben
\chi\t(12)=\chi\s\t(12) + \int d3d4\, \chi\s\t(13) f\Hxc\t(34) \chi\t(42).
\label{thGK}
\een
A simple approximation is then
the thermal adiabatic local
density approximation (thALDA), in which
the thermal XC kernel is approximated using the XC free energy density per particle
for a finite-temperature uniform gas, $a\xc^{\tau,{\rm unif}}$:
\ben
f\xc^{\tau,{\rm thALDA}}[\n](\br,\br',\omega) = 
\frac{d^2 a\xc^{\tau,{\rm unif}}(\n)}{d^2 \n}\Bigg|_{\n(\br)}\, \delta(\br-\br'),
\label{thALDA}
\een
which ignores its nonlocality in space and time, and could be used to generalize
ALDA calculations of excitations in metals and their surfaces\cite{Liebsch97}.

We next deduce the fluctuation-dissipation theorem for
MKS thermal DFT calculations. This allows us to connect the response function and the Coulomb interaction
through the dynamical structure factor\cite{LP77}.
In the MKS scheme, the XC contributions to the free energy are defined via
\begin{align}
A\t[n]&=T[n]+V\ee[n]+V[n]-\tau S[n]\\
&=T\s[n]+U[n]+V[n]-\tau S\s[n]+A\xc\t[n].
\end{align}
By subtraction,
\ben
A\t\xc[n]=T\t\xc[n]+U\xc\t[n]-\tau S\xc\t[n]
\een
where $T$ denotes kinetic, $U$ denotes potential, and $S$ entropic components.
Using many-body theory, the density-density response function determines the
potential contribution to correlation\cite{FW71,U11}, just as in the ground
state\cite{D12}:
\begin{align}
U\c\t&=&V\ee[\hat{\Gamma}^\tau[n]] - V\ee[\hat{\Gamma}\s^\tau[n]]~~~~~~~~~~~~\\
&=&-\int d\br \int d\br' \int_0^\infty \frac{d\omega}{2\pi}\coth{\bigg(\frac{\omega}{2\tau}\bigg)}  
\frac{\Im \Delta \chi\t(\br,\br',\omega)}{|\br-\br'|},
\label{Uct}
\end{align}
where $\Delta\chi\t=\chi\t-\chi\s\t$.
By introducing a coupling-constant $\lambda$
while keeping the density fixed, the thermal connection formula\cite{PB15}
yields
\begin{align}
&A\c\t[\n]=\notag\\
&\lim_{\tau''\to\infty}\frac{\tau}{2} \int_\tau^{\tau''} \frac{d\tau'}{\tau'^2}\, 
\int d\br\int d\br'\int \frac{d\omega}{2\pi}\coth{\bigg(\frac{\omega}{2\tau}\bigg)} \frac{\Im\Delta\chi^{\tau'}[\n\g](\br,\br',\omega)}
{|\br-\br'|},
\label{AcFD}
\end{align}
where the scaled density is $\n\g(\br) = \gamma^3 \n(\gamma\br)$ and $\gamma=\sqrt{\tau'/\tau}$.
This is exact, but only if the exact thermal XC kernel is used, as defined by
Eq. (\ref{kernel}).  If the kernel
is omitted, the result is the thermal random-phase approximation\cite{BWG05}.

Next, we discuss the many applications of Eq. (\ref{AcFD}).
There has been tremendous progress in implementing and testing the random phase approximation
for calculating the XC energy in ground-state calculations and such calculations, while
more expensive than standard DFT, are becoming routine\cite{EYF10,EF11,PRRS12}.  
Our results provide a thermal
generalization that could likewise be used to generate new thermal XC approximations
for equilibrium WDM calculations.
At finite temperature, the XC hole
fails to satisfy the simple sum rules\cite{PK98b} that have proven
so powerful in constructing ground-state
approximations\cite{BPW97}.   But our formula uses
instead the XC kernel. Inserting Eqs (\ref{thGK},\ref{thALDA}) into Eq. (\ref{AcFD}) yields
thALDA-RPA, a new approximation to the equilibrium correlation energy, that
can be applied to any system.
Another, simpler approximation is ALDA, in which
only the zero-temperature XC energy is used in the kernel.  Both can be
relatively easily evaluated for a uniform gas, and the resulting $a\xc\t(r\s)$
found from  Eq. (\ref{AcFD}) compared with an accurate parametrization\cite{KSDT14}.
Even in the uniform gas, thALDA is an approximation because both the $q$- and
$\omega$-dependence of the true $f\xc\t$
are missing; thus the efficacy of these approximations can be tested on the uniform case.

Next we discuss which known exact conditions on the zero-temperature kernel apply to the thermal kernel,
and which do not.  Because the equilibrium solution is a
minimum of the thermal free-energy functional,
the zero-force theorem\cite{U11} 
\ben
\int d^3 r\int d^3 r' \n\t(\br)\n\t(\br') f\xc\t(\br,\br',\omega)=0
\een
should be satisfied and
the kernel should be symmetric in its
spatial arguments. However, any simple formula for a one-electron system\cite{WYB12} is not
true at finite temperature, as the particle number is only an average in the grand canonical ensemble\cite{PK98b,CDK15}.

A last set of conditions is found by considering the coupling-constant
dependence in DFT.  A parameter $\lambda$ is introduced that multiplies the
electron-electron interaction, while keeping the density constant.  Because of
simple scaling relations, the $\lambda$-dependence can be shown to be
determined entirely by coordinate scaling of the density as in Eq. (\ref{AcFD}), i.e.,
determined by the functional itself, evaluated at different densities.
This is used in both ground-state DFT\cite{LP85} and in time-dependent
DFT\cite{HPB99}, and has been generalized to the thermal
case\cite{PPFS11,PB15}.  Although the thermal connection formula does not
require this relation for the response function, it is useful in many contexts.
From the Lehmann representation\cite{SL13} of $\chi\t$\cite{FW71}, we find
the $\lambda$-dependent response function satisfies:
\ben
\chi\tl[\n](\br,\br',\omega)=\lambda^4\, \chi^{\tau/\lambda^2}[n_{1/\lambda}](\lambda\br,\lambda\br',\omega/\lambda^2).
\label{chitl}
\een
Insertion into the definition of $f\xc$ yields:
\ben
f\xc\tl[\n](\br,\br',\omega)=\lambda^2\, f\xc^{\tau/\lambda^2}[n_{1/\lambda}](\lambda\br,\lambda\br',\omega/\lambda^2),
\label{fxctl}
\een
and the potential perturbation scales as:
\ben
\delta v\xc\tl[\n](\br,\omega)=\lambda^2\, \delta v\xc^{\tau/\lambda^2}[\n_{1/\lambda}](\lambda\br,\omega/\lambda^2).
\een
Insertion of the scaling relation for the
kernel into the thermal connection formula yields a more familiar
analog to the ground-state formula.

The exchange kernel must scale linearly with coupling constant, so Eq. (\ref{fxctl}) produces a rule
for scaling of the exchange kernel:
\ben
f\x\t[\n\g](\br,\br',\omega)=\gamma 
f\x^{\tau/\gamma^2}[n](\gamma\br,\gamma\br',\omega/\gamma^2).
\label{fxtg}
\een
Because the poles in $f\xc$ are $\lambda$-dependent, we expect pathologies similar to those in
zero-temperature TDDFT if 
the exact frequency-dependent $f\x\t$ is used in Eq. (\ref{AcFD})\cite{HB08}.
But adiabatic EXX (AEXX), not including frequency-dependence, produces a well-defined
approximation to the thermal free energy in which the kernel is non-local.  
This and the other proposed approximations above could prove useful
in WDM simulations when thermal XC effects are relevant (but see \cite{SPB15} for discussion of the subtleties involved in thermal XC approximations).

In conclusion, we have generalized the proofs and constructions of
TDDFT within the linear response formalism to thermal
ensembles, including those containing a finite number of degeneracies.
We have avoided ambiguities about the relative
perturbative and thermal equilibration time scales,
allowed for degenerate excited states more common in finite-temperature ensembles, avoided invoking boundary conditions and the requirement of 
Taylor expandability, and provided firm footing for  finite-temperature,
time-dependent KS-DFT in the linear response regime.
Definition of relevant KS quantities led to description of their properties under scaling.  
Further, we have shown that these quantities, in combination with
the thermal connection formula, produce new routes
to thermal DFT approximations for use in equilibrium MKS calculations.
Implementation and tests of these approximations is ongoing.

APJ acknowledges support from DE-FG02-97ER25308 and the University of California President's Postdoctoral Fellowship, PG from DE14-017426, and KB
from CHE-1464795 NSF.  Part of this work was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract DE-AC52-07NA27344.

\label{page:end}
\end{document}